# Seeing and Knowing in the Wild: Open-domain Visual Entity Recognition with Large-scale Knowledge Graphs via Contrastive Learning

### v1

@@ Introduction
(a)(b)Refer to caption Entity ID: Q160659 Name: London Eye Q: What is the wheel here? Entity: Q84 Name: London Q: What is the city in the picture? Entity: Q456960 Name: Westm. Bridge Q: What is the bridge in the front? Entity: Q41225 Name: Big Ben Q: What is clock tower on the left? Entity: Q62408 Name: Palace of Westm. Q: What is palace near the Big Ben? Refer to captionText SpaceImage SpaceSemantic SpaceLondonPalace of Westm.Big BenLondon Eye Ferris Wheel Wheel Tourist Attract. Westm. BridgeRoad. BridgeBridgeP276P276P276P31P31P31P31P31P361P279P279 P276 = Location P31 = Instance of P361 = Part of P279 = Sub Class of Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption The Palace of Westminster is the meeting place of the Parliament of the United King… Big Ben is the nickname for the Great Bell of the Great Clock of Westm… The London Eye, originally the Millennium Wheel, is a cantilevered observation wheel on… Interclass Dist.Intraclass Dist.Modality Gap
@@
    We present a dual-encoder approach for open-domain visual entity recognition task that leverages external knowledge from both Wikidata and Wikipedia to enable a better representation of images and text descriptions.
@@
Contrastive loss is a widely-used objective in representation learning, particularly in contexts such as metric learning, image-text alignment, and self-supervised learning. Given two sets of embeddings $A=\{a_{i}\}_{i=1}^{N}$ and $B=\{b_{i}\}_{i=1}^{N}$, where $(a_{i},b_{i})$ is a positive (matching) pair and $\forall(a_{i},b_{j}),j\neq i$, is a negative pair, we define the contrastive loss as:
@@
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $\tau$ is a temperature parameter controlling the softness of the similarity distribution, and $N$ is the set size. Common similarity functions include cosine similarity and Euclidean distance-based similarity. Specifically, the symmetric contrastive loss can be defined as:
$\ell_{\text{sym}}(A,B)=\frac{1}{2}\left(\ell\left(A,B\right),\ell\left(B,A\right)\right)\text{,}$ (2)
which encourages mutual alignment and it is widely used in various models like CLIP.
@@
Wikidata is a collaboratively-maintained, multilingual knowledge graph that can be formalised as the directed graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is the set of entities (called items and uniquely referenced by “Q-identifiers” (QIDs), e.g., Q42 for Douglas Adams), $\mathcal{R}$ is the set of relation types (called properties and referenced by ”P-identifiers“ (PIDs), e.g., P31 for instance of), and $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ is the set of knowledge triples $(e_{1},r,e_{2})$ stating that subject entity $e_{1}$ is linked to object entity $e_{2}$ via property $r$. As of 2025, Wikidata KG contains more than 100 millions entities, 12,000 properties, and 10 billions of triples and is growing daily through human and bot contributions.
@@
Early approaches (Farhadi et al. 2009; Lampert, Nickisch, and Harmeling 2009) introduced human-defined attributes as intermediate semantic representations for recognition of unseen classes. Later methods moved toward unsupervised semantic embeddings derived from text corpora (Socher et al. 2013; Frome et al. 2013; Norouzi et al. 2014). These dual encoder or projection-based approaches have proven that aligning images with distributed text representations enables zero-shot recognition at scale. Vision-language pretraining has greatly advanced this paradigm. A well-known example is CLIP (Radford et al. 2021) that is trained on 400 million image-text pairs to learn a joint embedding space for images and natural language. By encoding an image and a candidate label (or description) into the same space, CLIP can directly measure their compatibility, essentially performing zero-shot classification by ranking labels. CLIP has wide applications in visual inspection (Jeong et al. 2023; Sadikaj et al. 2025) and robotics (Zhou et al. 2023, 2024a; Yao et al. 2025).
Another line of zero-shot research uses large-scale generative models to bridge the gap to unseen classes. Instead of directly mapping images to text semantics, these generative approaches output the entity names based on given images and then link the entity names to the categories (two-step approach). For instance, Hu et al. (2023) uses PaLI and BLIP-v2 (Li et al. 2023a) on the OVEN benchmark to make the model predict the entity name and then use the BM25 search algorithm to find the closed label in the label space of Wikidata ids. Instead of generating entity names, GER-ALD (Caron et al. 2024) generates the semantic and discriminative “code” to identify entities. Auto-VER (Xiao et al. 2024) combines contrastive learning with generative models to enhance the ability to distinguish similar entities within a vast label space. Unlike dual-encoder approaches, two-step generative methods output text conditioned on images and then match them to labels in the label space. This introduces semantic ambiguity, as generated outputs may be correct lexically but vague or incomplete semantically.
@@
Many studies have explored incorporating knowledge into vision-and-language tasks (Monka, Halilaj, and Rettinger 2022), including visual question answering (Chang et al. 2022; Chen et al. 2021) and entity-aware image captioning (Biten et al. 2019). For the visual entity recognition task, Wang, Ye, and Gupta (2018) and Kampffmeyer et al. (2019) use semantic embeddings and categorical relationships from KGs through graph convolution networks for zero-shot prediction. Li et al. (2023b) augments few-shot image recognition by introducing auxiliary semantic prior knowledge and propagating knowledge among categories via a semantic-visual mapping. KG-NN (Monka et al. 2021) combines prior knowledge encoded in KGs with visual representations to enhance generalization under distribution shifts, and KGV (Zhou et al. 2024b) demonstrates that the integration of richer multi-modal priors can further improve performance. Instead of relying on small domain-specific knowledge graphs, our approach is the first to leverage a large-scale knowledge graph for open-domain visual entity recognition, demonstrating strong scalability potential.
@@
Our goal is to integrate the rich structured knowledge from Wikidata into the training pipeline. To achieve this, we align knowledge graph embeddings (KGEs), image embeddings, and text embeddings within a shared latent space, enabling the structured knowledge captured by the KGEs to be effectively injected into the latent space’s representation. A framework overview can be seen in Figure 2.
@@
We employ a dual-encoders approach to align the semantic representation of the input image and text query with that of the corresponding entity, which includes its lead image(s) and textual description. We structure our discussion into two parts: (1) encoding the input image and text query, and (2) encoding the multi-modal information of entities from the knowledge graph.
@@
with parameters $\mathbf{\gamma}$ to to fuse the image embedding $f_{\mathbf{\theta}}(x^{p})$ and text embedding $f_{\mathbf{\lambda}}(x^{t})$. The embedding, combining the information from the input image and text query, can be written as
@@
For given input batch $\{(x^{p}_{i},x^{t}_{i})\}_{i=1}^{N_{b}}$, where $x^{p}_{i}$ and $x^{t}_{i}$ are the $i$-th input images and its corresponding text query, $N_{b}$ is the batch size, the corresponding ground truth label set is denoted as $\{e_{i}\}_{i=1}^{N_{b}}$. We define three losses, namely alignment loss, proxy loss, and knowledge graph loss for the training process. Figure 3 shows the learning visualization.
@@
$\mathcal{T}_{\text{(h,r,t)}}=\{(h^{\prime},r,t)\mid h^{\prime}\in\mathcal{E}\setminus h\}\cup\{(h,r,t^{\prime})\mid t^{\prime}\in\mathcal{E}\setminus h\}\text{.}$ (8)
@@
In our experiments, we investigate whether incorporating knowledge from large-scale KGs and KBs can enhance open-domain visual entity recognition. We analyze the impact of different types of prior knowledge, fusion methods, and KGE techniques. In addition, ablation studies on various hyperparameter values are also provided.
@@
Git-Large (Caron et al. 2024) CVPR2024 0.4 WebLI 17.6 4.3 4.0
GER-ALD (Caron et al. 2024) CVPR2024 0.4 LAION 29.1 17.7 22.7
@@
The OVEN dataset contains 6,063,945 training samples, which combines 14 image recognition datasets. Models are evaluated on the test splits, and performance is reported using the harmonic mean (HM) of top-1 accuracy for two types of entities: seen entities, which appear in the OVEN training set, and unseen entities, which are not present during training. Specifically, the test split includes 15,888 entities (8,355 seen and 7,533 unseen). Our evaluation is conducted solely on the OVEN benchmark, comprising 14 diverse image recognition datasets, thus indicating the generalizability of our approach. The evaluation metric is top-1 accuracy for seen and unseen entities, and the harmonic mean (HM) of them.
@@
We extract a subgraph from Wikidata containing 32,122 entities and 501 relation types. For training, we use the AdamW optimizer with a learning rate of 0.001, batch size of 4096, and weight decay of 0.0001. Our approach includes three model variants based on different OpenCLIP image encoders: ViT-L/14, ViT-H/14, and ViT-bigG/14. The temperature $\tau$ for contrastive learning is fixed at 0.07 during fine-tuning. The hyperparameters $\beta_{1}$ and $\beta_{2}$ are both set to 1. We fuse information from the input image and text query (Equation (3)), using a simple addition operation. More details are provided in the Appendix.
@@
Dual encoders approaches consist of two encoders, each encoding a dedicated modality (e.g., images and text) into the same latent space. The visual recognition is conducted by searching the nearest neighbors of prototypes in the latent space for a given input image. Representative models include CLIP, CLIPfusion, and CLIP2CLIP.
@@
Another approach uses auto-regressive models to generate captions for the given images. Representative vision-language models (VLMs) such as PaLI, GiT-Large, GER-ALD, Auto-VER follow this paradigm. These models generate a descriptive caption or name for the visual entity and then match it to the closest label in the predefined label space for final prediction.
@@
Here, we analyse the impact of model size, prior knowledge types, fusion strategy of the text query and input image, various KGE methods, and hyperparameters - $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and temperature $\tau$. The detailed setting can be found in the Appendix section experiment setting.
@@
We highlight that KnowCoL-L, CLIP, CLIPFusion, and CLIP2CLIP use the same backbone ViT-L-14, but KnowCoL-L significantly outperforms all. While CLIPFusion and CLIP2CLIP achieve harmonic means of 8.4% and 11.4% respectively, KnowCoL-L reaches 31.5%, showing over 3 times improvement. This demonstrates that external knowledge significantly improves zero-shot generalization without increasing model size.
@@
Table 3: Impact of different types of knowledge. We investigate the impact of structural knowledge, namely the sample images in the KB and the hierarchical relations in the KG.
Table 3 demonstrates the importance of different knowledge types in the KnowCoL-L model. The complete KnowCoL-L model achieves the best score of 31.5% HM. Removing lead images (w/o Lead images) reduces performance to 32.5% (seen), 27.9% (unseen), and 30.3% (HM), highlighting the role of visual context. Omitting hierarchical KG knowledge of ‘Instance of’, ‘subclass of’, and ‘parent taxon’ further lowers scores to 32.1% (seen), 25.3% (unseen), and 28.3% (HM), significantly impairing unseen entity recognition.
@@
As Table 4 shows, the default Addition method provides balanced performance, achieving the best harmonic mean 31.5%. By using the multi-layer perceptron (MLP) method, the performance for seen entities improves up to 37.5%, while for unseen entities decreases to 26.8%. Using transformer encoder (TE) methods with or without local image patch embeddings, significantly boost the accuracy on seen entities (54.4% and 56.7%), but degrade unseen entity recognition (16.6% and 17.9%). The trend shows that more complex fusion methods, containing more layers or a more complex structure, lead to overfitting on seen entities and sacrificing zero-shot generalization to unseen entities. Additionally, by introducing local image patch tokens (TE + IP), performance on both seen (56.7%) and unseen entities (17.9%) improves compared to the TE-only method. This improvement likely occurs because local image patch tokens provide fine-grained visual details and enhance recognition ability.
@@
Table 5 compares various knowledge graph embedding methods and distance measures. Our default choice TransE (Bordes et al. 2013) with cosine similarity, achieves the best overall performance. This setup aligns well with CLIP, which also uses cosine similarity in its contrastive learning. When replacing the cosine similarity with Euclidean distance, the HM drops from 31.5% to 31.0%. TransH (Wang et al. 2014) introduces more representational flexibility by allowing entity embeddings to vary across different relations. This additional modeling capacity slightly improved seen performance to 34.4% but comes at the cost of lower unseen accuracy to 28.6%. The DistMult method (Yang et al. 2015) shows the lowest performance, suggesting it is less effective for integrating structural knowledge from the KG into the visual-textual space.
@@
As Figure 4 shows, we conduct experiments varying the hyperparameters $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and contrastive learning temperature $\tau$. Figure 4a shows that increasing the weight of the proxy loss $\beta_{1}$, aligning node embeddings with multi-modal prior knowledge, enhances model performance. Figure 4b demonstrates that increasing $\beta_{2}$, the weight of KE loss incorporating structured prior knowledge from Wikidata, also improves performance. Additionally, we identify that the larger model (KnowCoL-H) with more prior knowledge can benefit more from the inclusion of structured prior knowledge. Figure 4c illustrates that larger models require higher dimensional latent spaces to represent richer features. The temperature 0.07 performs best in our task, as shown in Figure 4d.
Conclusion and Future Works
We propose KnowCoL, a novel approach to improve open-domain visual entity recognition by incorporating multi-modal prior knowledge, namely structured knowledge, textual and visual priors. KnowCoL significantly enhances visual entity recognition performance, especially for zero-shot recognition of unseen entities. Future work includes exploring non-Euclidean spaces like hyperbolic and hyperspherical spaces to better capture the hierarchical relationships and semantic structures inherent in knowledge graphs. Investigating these non-Euclidean geometries could further improve representation learning, potentially leading to more accurate and efficient entity recognition and generalization.
Acknowledgement
This work was partially funded by the European Union’s Horizon RIA research and innovation programme under grant agreement No. 101092908 (SMARTEDGE). The authors also thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Hongkuan , Yuqicheng Zhu, and Jingcheng Wu.

### v2

@@ Introduction
(a)(b)Refer to caption Entity ID: Q41225 Name: Big Ben Q: What is the clock tower? Entity: Q84 Name: London Q: What is the city in the picture? Entity: Q18536 Name: The Shard Q: What is the glass building on the right? Entity: Q160659 Name: London Eye Q: What is the wheel here? Entity: Q62408 Name: Palace of Westm. Q: What is palace near the Big Ben? Refer to captionText SpaceImage SpaceSemantic SpaceLondonPalace of Westm.Big BenLondon Eye Ferris Wheel Wheel Tourist Attract. Westm. BridgeRoad. BridgeBridgeP276P276P276P31P31P31P31P31P361P279P279 P276 = Location P31 = Instance of P361 = Part of P279 = Sub Class of Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption The Palace of Westminster is the meeting place of the Parliament of the United King… Big Ben is the nickname for the Great Bell of the Great Clock of Westm… The London Eye, originally the Millennium Wheel, is a cantilevered observation wheel on… Interclass Dist.Intraclass Dist.Modality Gap
@@
    We present a dual-encoder approach for the open-domain visual entity recognition task that leverages external knowledge from both Wikidata and Wikipedia to enable a better representation of images and text descriptions.
@@
Contrastive loss is a widely used objective in representation learning, particularly in contexts such as metric learning, image-text alignment, and self-supervised learning. Given two sets of embeddings $A=\{a_{i}\}_{i=1}^{N}$ and $B=\{b_{i}\}_{i=1}^{N}$, where $(a_{i},b_{i})$ is a positive (matching) pair and $\forall(a_{i},b_{j}),j\neq i$, is a negative pair, we define the contrastive loss as:
@@
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $\tau$ is a temperature parameter controlling the softness of the similarity distribution, and $N$ is the set size. Common similarity functions include cosine similarity and Euclidean distance-based similarity. Specifically, the symmetric contrastive loss can be defined as,
$\ell_{\text{sym}}(A,B)=\frac{1}{2}\left(\ell\left(A,B\right)+\ell\left(B,A\right)\right)\text{,}$ (2)
which encourages mutual alignment, and it is widely used in various models like CLIP.
@@
Wikidata is a collaboratively-maintained, multilingual knowledge graph that can be formalised as the directed graph $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is the set of entities (called items and uniquely referenced by “Q-identifiers” (QIDs), e.g., Q42 for Douglas Adams), $\mathcal{R}$ is the set of relation types (called properties and referenced by ”P-identifiers“ (PIDs), e.g., P31 for instance of), and $\mathcal{T}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ is the set of knowledge triples $(e_{1},r,e_{2})$ stating that subject entity $e_{1}$ is linked to object entity $e_{2}$ via property $r$. As of 2025, Wikidata KG contains more than 100 million entities, 12,000 properties, and 10 billion triples and is growing daily through human and bot contributions.
@@
Early approaches (Farhadi et al. 2009; Lampert, Nickisch, and Harmeling 2009) introduced human-defined attributes as intermediate semantic representations for recognition of unseen classes. Later methods moved toward unsupervised semantic embeddings derived from text corpora (Socher et al. 2013; Frome et al. 2013; Norouzi et al. 2014). These dual-encoder or projection-based approaches have proven that aligning images with distributed text representations enables zero-shot recognition at scale. Vision-language pretraining has greatly advanced this paradigm. A well-known example is CLIP (Radford et al. 2021), which is trained on 400 million image-text pairs to learn a joint embedding space for images and natural language. By encoding an image and a candidate label (or description) into the same space, CLIP can directly measure their compatibility, essentially performing zero-shot classification by ranking labels. CLIP has wide applications in visual inspection (Jeong et al. 2023; Sadikaj et al. 2025) and robotics (Zhou et al. 2023, 2024a; Yao et al. 2025).
Another line of zero-shot research uses large-scale generative models to bridge the gap to unseen classes. Instead of directly mapping images to text semantics, these generative approaches output the entity names based on given images and then link the entity names to the categories (two-step approach). For instance, Hu et al. (2023) use PaLI and BLIP-v2 (Li et al. 2023a) on the OVEN benchmark to make the model predict the entity name and then uses the BM25 search algorithm to find the closest label in the label space of Wikidata ids. Instead of generating entity names, GER-ALD (Caron et al. 2024) generates the semantic and discriminative “code” to identify entities. Auto-VER (Xiao et al. 2024) combines contrastive learning with generative models to enhance the ability to distinguish similar entities within a vast label space. Unlike dual-encoder approaches, two-step generative methods output text conditioned on images and then match them to labels in the label space. This introduces semantic ambiguity, as generated outputs may be correct lexically but vague or incomplete semantically.
@@
Many studies have explored incorporating knowledge into vision-and-language tasks (Monka, Halilaj, and Rettinger 2022), including visual question answering (Chang et al. 2022; Chen et al. 2021) and entity-aware image captioning (Biten et al. 2019). For visual entity recognition task, Wang, Ye, and Gupta (2018) and Kampffmeyer et al. (2019) use semantic embeddings and categorical relationships from KGs through graph convolution networks for zero-shot prediction. Li et al. (2023b) augments few-shot image recognition by introducing auxiliary semantic prior knowledge and propagating knowledge among categories via a semantic-visual mapping. KG-NN (Monka et al. 2021) combines prior knowledge encoded in KGs with visual representations to enhance generalization under distribution shifts, and KGV (Zhou et al. 2024b) demonstrates that the integration of richer multi-modal priors can further improve performance. Instead of relying on small domain-specific knowledge graphs, our approach is the first to leverage a large-scale knowledge graph for open-domain visual entity recognition, demonstrating strong scalability potential.
@@
Our goal is to integrate the rich, structured knowledge from Wikidata into the training pipeline. To achieve this, we align knowledge graph embeddings (KGEs), image embeddings, and text embeddings within a shared latent space, enabling the structured knowledge captured by the KGEs to be effectively injected into the latent space’s representation. A framework overview can be seen in Figure 2.
@@
We employ a dual-encoder approach to align the semantic representation of the input image and text query with that of the corresponding entity, which includes its lead image(s) and textual description. We structure our discussion into two parts: (1) encoding the input image and text query, and (2) encoding the multi-modal information of entities from the knowledge graph.
@@
with parameters $\mathbf{\gamma}$ to fuse the image embedding $f_{\mathbf{\theta}}(x^{p})$ and text embedding $f_{\mathbf{\lambda}}(x^{t})$. The embedding, combining the information from the input image and text query, can be written as
@@
For given input batch $\{(x^{p}_{i},x^{t}_{i})\}_{i=1}^{N_{b}}$, where $x^{p}_{i}$ and $x^{t}_{i}$ are the $i$-th input images and its corresponding text query, $N_{b}$ is the batch size, the corresponding ground truth label set is denoted as $\{e_{i}\}_{i=1}^{N_{b}}$. We define three losses, namely alignment loss, proxy loss, and knowledge graph loss, for the training process. Figure 3 shows the learning visualization.
@@
$\mathcal{T}_{\text{(h,r,t)}}=\{(h^{\prime},r,t)\mid h^{\prime}\in\mathcal{E}\setminus\{h\}\}\cup\{(h,r,t^{\prime})\mid t^{\prime}\in\mathcal{E}\setminus\{t\}\}\text{.}$ (8)
@@
In our experiments, we investigate whether incorporating knowledge from large-scale KGs and Knowledge Base (KB) can enhance open-domain visual entity recognition. We analyze the impact of different types of prior knowledge, fusion methods, and KGE techniques. In addition, ablation studies on various hyperparameter values are also provided.
@@
GIT-Large (Caron et al. 2024) CVPR2024 0.4 WebLI 17.6 4.3 7.0
GER-ALD (Caron et al. 2024) CVPR2024 0.4 LAION 29.1 16.3 20.9
@@
The OVEN dataset contains 6,063,945 training samples, which combines 14 image recognition datasets. Models are evaluated on the test splits, and performance is reported using the harmonic mean (HM) of top-1 accuracy for two types of entities: seen entities, which appear in the OVEN training set, and unseen entities, which are not present during training. Specifically, the test split includes 15,888 entities (8,355 seen and 7,533 unseen). Our evaluation is conducted solely on the OVEN benchmark, comprising 14 diverse image recognition datasets, thus indicating the generalizability of our approach. The evaluation metric is top-1 accuracy for seen and unseen entities and the HM of them.
@@
We extract a subgraph from Wikidata containing 32,122 entities and 501 relation types. For training, we use the AdamW optimizer with a learning rate of 0.001, batch size of 4096, and weight decay of 0.0001. Our approach includes three model variants based on different OpenCLIP image encoders: ViT-L/14, ViT-H/14, and ViT-bigG/14. The temperature $\tau$ for contrastive learning is fixed at 0.07 during fine-tuning. The hyperparameters $\beta_{1}$ and $\beta_{2}$ are both set to 1. We fuse information from the input image and text query (Equation (3)), using a simple addition operation. More details are in the Appendix.
@@
Dual-encoder approaches consist of two encoders, each encoding a dedicated modality (e.g., images and text) into the same latent space. The visual recognition is conducted by searching the nearest neighbors of prototypes in the latent space for a given input image. Representative models include CLIP, CLIPfusion, and CLIP2CLIP.
@@
Another approach uses auto-regressive models to generate captions for the given images. Representative vision-language models (VLMs) such as PaLI, GIT-Large, GER-ALD, and Auto-VER follow this paradigm. These models generate a descriptive caption or name for the visual entity and then match it to the closest label in the predefined label space for final prediction.
@@
Here, we analyse the impact of model size, prior knowledge types, fusion strategy of the text query and input image, various KGE methods, and hyperparameters - $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and temperature $\tau$. The detailed setting can be found in the Appendix section, experiment setting.
@@
We highlight that KnowCoL-L, CLIP, CLIPFusion, and CLIP2CLIP use the same backbone ViT-L-14, but KnowCoL-L significantly outperforms all. While CLIPFusion and CLIP2CLIP achieve harmonic means of 8.4% and 11.4%, respectively, KnowCoL-L reaches 31.5%, showing about 2.8 times improvement. This demonstrates that external knowledge significantly improves zero-shot generalization without increasing model size.
@@
Table 3: Impact of different types of knowledge. We investigate the impact of structural knowledge, namely the Lead images in the KB and the hierarchical relations in the KG
Table 3 demonstrates the importance of different knowledge types in the KnowCoL-L model. The complete KnowCoL-L model achieves the best score of 31.5% HM. Removing lead images (w/o Lead images) reduces performance to 32.5% (seen), 27.9% (unseen), and 30.0% (HM), highlighting the role of visual context. Omitting hierarchical KG knowledge of ‘Instance of’, ‘subclass of’, and ‘parent taxon’ further lowers scores to 32.1% (seen), 25.3% (unseen), and 28.3% (HM), significantly impairing unseen entity recognition.
@@
As Table 4 shows, the default Addition method provides balanced performance, achieving the best harmonic mean 31.5%. By using the multi-layer perceptron (MLP) method, the performance for seen entities improves up to 37.5%, while for unseen entities decreases to 26.8%. Using transformer encoder (TE) methods with or without local image patch embeddings significantly boost the accuracy on seen entities (54.4% and 56.7%), but degrade unseen entity recognition (16.6% and 17.9%). The trend shows that more complex fusion methods, containing more layers or a more complex structure, lead to overfitting on seen entities and sacrificing zero-shot generalization to unseen entities. Additionally, by introducing local image patch tokens (TE + IP), performance on both seen (56.7%) and unseen entities (17.9%) improves compared to the TE-only method. This improvement likely occurs because local image patch tokens provide fine-grained visual details and enhance recognition ability.
@@
Table 5 compares various knowledge graph embedding methods and distance measures. Our default choice, TransE (Bordes et al. 2013) with cosine similarity, achieves the best overall performance. This setup aligns well with CLIP, which also uses cosine similarity in its contrastive learning. When replacing the cosine similarity with Euclidean distance, the HM drops from 31.5% to 31.0%. TransH (Wang et al. 2014) introduces more representational flexibility by allowing entity embeddings to vary across different relations. This additional modeling capacity slightly improved the seen performance to 34.4% but comes at the cost of lower unseen accuracy 28.6%. The DistMult method (Yang et al. 2015) shows the lowest performance, suggesting it is less effective for integrating structural knowledge from the KG into the visual-textual space.
@@
As Figure 4 shows, we conduct experiments varying the hyperparameters $\beta_{1}$, $\beta_{2}$, latent space dimension $d_{e}$, and contrastive learning temperature $\tau$. Figure 4a shows that increasing the weight of the proxy loss $\beta_{1}$, aligning node embeddings with multi-modal prior knowledge, enhances model performance. Figure 4b demonstrates that increasing $\beta_{2}$, the weight of KE loss incorporating structured prior knowledge from Wikidata, also improves performance. Additionally, we identify that the larger model (KnowCoL-H) with more prior knowledge can benefit more from the inclusion of structured prior knowledge. Figure 4c illustrates that larger models require higher-dimensional latent spaces to represent richer features. The temperature 0.07 performs best in our task, as shown in Figure 4d.
Conclusion and Future Work
We propose KnowCoL, a novel approach to improve open-domain visual entity recognition by incorporating multi-modal prior knowledge, namely structured knowledge, textual, and visual priors. KnowCoL significantly enhances visual entity recognition performance, especially for zero-shot recognition of unseen entities. Future work includes exploring non-Euclidean spaces like hyperbolic and hyperspherical spaces to better capture the hierarchical relationships and semantic structures inherent in knowledge graphs.
Acknowledgments
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Hongkuan Zhou, Yuqicheng Zhu, and Jingcheng Wu. This work was partially funded by the European Union’s Horizon RIA research and innovation programme under grant agreement No. 101092908 (SMARTEDGE). Jingcheng Wu has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - SFB 1574 - Project number 471687386. The authors also gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research, and the Arts of Baden-Württemberg, as part of the National High-Performance Computing (NHR) joint funding program (https://www.nhr-verein.de/en/our-partners). HoreKa is partly funded by the German Research Foundation (DFG).